# A novel characterization of structures in smooth regression curves: from a viewpoint of persistent homology

### v3

@@ Abstract
We characterize structures such as monotonicity, convexity, and modality in smooth regression curves using persistent homology. Persistent homology is a key tool in topological data analysis that detects higher dimensional topological features such as connected components and holes (cycles or loops) in the data. In other words, persistent homology is a multiscale version of homology that characterizes sets based on the connected components and holes. We use super-level sets of functions to extract geometric features via persistent homology. In particular, we explore structures in regression curves via the persistent homology of super-level sets of a function, where the function of interest is - the first derivative of the regression function.
In the course of this study, we extend an existing procedure of estimating the persistent homology for the first derivative of a regression function and establish its consistency. Moreover, as an application of the proposed methodology, we demonstrate that the persistent homology of the derivative of a function can reveal hidden structures in the function that are not visible from the persistent homology of the function itself. In particular, we characterize structures such as monotonicity, convexity, and modality, and propose a measure of statistical significance to infer these structures in practice. Finally, we conduct an empirical study to implement the proposed methodology on simulated and real data sets and compare the derived results with an existing methodology.
@@ 1 Introduction
There have been various attempts to explore the shape of the regression function from the geometric perspective, for instance; [28] proposed a rank correlation coefficient to compare the shape of regression functions. In a different line of approach, motivated by the scale-space view from computer vision literature, [6] proposed a graphical device called SiZer to explore structures, such as peaks and valleys, in regression curves. SiZer analysis has been used in comparing two or more regression curves (see, e.g., [33], [32]). [18] proposed a graphical device, based on the derivative and inverse of the derivative of the regression function, to compare the two regression functions up to shift in the horizontal and/or vertical axis. [20] compared the time-varying quantile regression curves up to the horizontal shift.
@@
We analyze super-level sets of a function from a topological perspective. In particular, we take the function to be analyzed as the first derivative of the regression function and then calculate the persistent homology of its super-level sets. Since the regression function is usually not known in practice, we estimate the persistent homology using an estimate of the derivative of the regression function. In this paper, we take the derivative of the Nadaraya-Watson estimator as an estimator of the derivative of the regression function. The validity of such estimators of persistent homology is due to the well-known stability theorem (see [11]). The stability theorem provides an upper bound of the bottleneck distance between true persistent homology and the estimated persistent homology in terms of the sup-norm distance between the underlying function and the corresponding estimator. Thus, if the estimator of the underlying function is uniformly convergent, then the corresponding persistent homology is consistent with respect to the bottleneck metric.
@@ 1.1 Our contribution
The persistent homology of a function does not reveal local geometric features of its own accord. It only reveals global topological features such as connected components and holes in higher dimensions. It is worth mentioning here that the use of persistent homology to study the geometric structures of 1-manifolds is mainly limited to the clustering structures. The use of persistent homology to infer local geometric structures is not so straightforward, and such procedures have been less developed in the literature. In this work, we study the local geometric features such as monotonicity, convexity, and modality of smooth regression functions not by studying their persistent homology but by exploring the persistent homology of their derivative. The methodology adopted here appears analogous to the procedure proposed by [1] to estimate the persistent homology of the derivative of the regression function, though they are entirely different too. In the present work, the goal is not only to recover the persistent homology of the true function but also to infer local geometric features of the underlying function using persistent homology. In contrast, the goal in [1] is to devise a new procedure to infer the persistent homology of density and regression functions. [1] established consistency of the proposed estimator of persistent homology for density and regression functions using their kernel estimators. In a complementary fashion, we extend those results to the first derivative of the regression function and establish consistency of the estimator of persistent homology up to the precision $5\epsilon$.
As an application of the proposed procedure, we demonstrate that it can be used to study the shape of regression functions. In particular, we characterize structures such as monotonicity, convexity, and modality in smooth regression curves and propose a measure of the statistical significance of these structures observed in the data. Moreover, one can adopt the proposed procedure to compare or cluster analysis of regression curves based on these structures. One of the real-life applications of our work is that given the regression curves of population growth of different subcontinents, one can investigate the shape of the first derivative of regression functions to compare the population growth patterns over time.
@@ 1.2 Mathematical challenges
Most of the techniques in the literature first obtain an estimator of the underlying function converging uniformly to its population counterpart and then use this estimator to estimate the persistent homology. As mentioned previously such procedures are supported by the stability results for persistent homology (see, Theorem 2.1).
@@ 1.3 Outline
The rest of the article is organized as follows. Section 2 gives the necessary topological background to elaborate the notion of persistent homology. Section 3 formulates the exact problem and the main results of the article. Section 5 presents some applications of the proposed methodology. Section 7 consists of some concluding remarks. Finally, proofs of the main theoretical results are supplied in the appendix.
@@ 2 Preliminaries
Persistent homology (see, e.g., [21], [25], [37]) is a prominent tool in topological data analysis (TDA) that provides a multi-scale summary of topological features present in the data. Persistent homology is a multiscale extension of homology which is one of the refined topological invariants from algebraic topology (see; e.g., [29], [27]). Heuristically, homology characterizes sets based on the number of connected components and holes using some algebraic machinery. In what follows, we are concerned with the homology of simplicial complexes referred to as simplicial homology. Simplicial complexes are combinatorial representations of topological spaces made up of points, lines, edges, triangles, and their higher dimensional analogs. In the following subsections, we briefly describe the necessary background and definitions related to the algebraic topology.
@@ 2.1 Simplicial homology
In this subsection, we define the procedure to compute the homology of simplicial complexes. Note that given a finite set of points in a Euclidean space, one can geometrically define simplicial complexes (see, e.g., [29], [27]). However, in TDA, data points may not be embedded into the Euclidean spaces, therefore one works with the abstract simplicial complexes. In what follows, abstract simplicial complexes are referred to as simplicial complexes.
@@
Let $\mathcal{X}=\{x_{0},\ldots,x_{n}\}$ be a finite set, and $\Delta$ be a collection of non-empty subsets in $\mathcal{X}$. Then $\Delta$ is said to be an abstract simplicial complex if it satisfies the following properties:
@@
The set $\mathcal{X}$ is also called the vertex set of $\Delta$ and the elements of $\mathcal{X}$ are referred to as vertices. Note that an abstract simplicial complex is completely determined by its set of vertices. For an integer $k\geq 0$, any subset $\sigma\in\Delta$ with $k+1$ vertices is said to be a k-simplex or a k-dimensional simplex. Any subset $\tau$ of a k-simplex $\sigma$ with k vertices is called a face of $\sigma$. The dimension of $\Delta$ denoted as dim($\Delta$) is defined to be the maximum dimension of simplices in $\Delta$. That is, dim(Δ) max $\text{dim}(\Delta)\triangleq\text{ max }\{\text{dim}(\sigma):\sigma\in\Delta\}$.
Now, we define the following algebraic structure on simplices in $\Delta$ to define holes or cycles in different dimensions of a topological space. First, we need to assign orientations; an ordering on vertices of simplices in $\Delta$. Let $\sigma=\{x_{i_{0}},\ldots,x_{i_{k}}\}$ be a k-simplex in $\Delta$, where $i_{0},\ldots,i_{k}$ is a permutation of $0,1,\ldots,k$. We say two ordering of its vertex set is equivalent if they differ from one another by an even permutation. Thus, orderings of the vertices of $\sigma$ fall into two equivalent classes. Each of these classes is referred to as an orientation of $\sigma$. A 0-simplex has only one orientation. For any $k\geq 0$, we denote an oriented simplex as $[\sigma]=[x_{i_{0}},\ldots,x_{i_{k}}$].
We denote the set of all k-simplices in $\Delta$ as $\Delta_{k}$. A k-chain on $\Delta$ is defined to be a formal linear combination of k-simplices in $\Delta_{k}$. Let $n_{k}=|\Delta_{k}|$, and $\gamma$ denotes a k-chain on $\Delta$, that is $\gamma\triangleq\sum\limits_{i=1}^{n_{k}}\alpha_{i}\sigma_{i}$, where the coefficients $\alpha_{i}$’s belong to some field $\mathbb{F}$. The space of k-chains is a vector space over the field $\mathbb{F}$, and defined as
@@
Now, we need to map the simplices in $\Delta_{k}$ into a basis of $C_{k}(\mathcal{X},\mathds{F})$ such that the map preserves orientations of simplices in $\Delta_{k}$. Let $\Delta^{\pi}_{k}$ denote the set containing all the simplices in $\Delta_{k}$ in all possible orientations. Any oriented simplex $[\sigma]$ in $\Delta^{\pi}_{k}$ is denoted as $[x_{\pi(0)},\ldots,x_{\pi(k)}]$, where $\pi=\{\pi(0),\ldots,\pi(k)\}$ is a permutation of $0,\ldots,k$. Then, for any simplex $\sigma_{i}\in\Delta_{k}$, define the map $T_{k}:\Delta^{\pi}_{k}\xrightarrow{}C_{k}(\mathcal{X},\mathds{F})$ as follows:
@@
where $\sigma$ is face of $\sigma_{i}$. The subspaces $B_{k}(\Delta)\triangleq\text{Im }\partial_{k+1}\text{ and }Z_{k}(\Delta)\triangleq\text{ ker }\partial_{k}$ are the set of k-boundaries and the set of k-cycles, respectively. The $k^{th}$ homology of $\Delta$ is defined to be the quotient vector space given by
@@ 2.2 Persistent homology
Let $\mathcal{X}$ be a space, then a filtration of $\mathcal{X}$ is a nested family of spaces $\{\mathcal{X}_{t}:t\in\mathbb{R}\}$ such that $\mathcal{X}=\bigcup_{t\in\mathbb{R}}\mathcal{X}_{t}$. There are two types of filtrations often used in TDA. First, filtrations built on top of a point cloud, for example, the families of Rips or $\check{C}$ech complex indexed over a set of non-negative real numbers. Second, filtrations obtained through level sets of a function $f:\mathcal{X}\xrightarrow{}\mathbb{R}$, for example, one can take $\mathcal{X}_{t}=\{x\in\mathcal{X}:f(x)\geq t\}$ to form a filtration of $\mathcal{X}$, such filtration is referred to as super-level set filtration of $f$. In practice, the index parameter $t$ is called the scale parameter. The homology of the spaces $\mathcal{X}_{t}$ varies with the scale parameter t. For example, as t increases, new connected components may be born, or existing connected components may disappear. Similarly, new holes or cycles are formed or filled up as t increases. Thus, as one moves through the filtration new features are born and existing features die. The birth time of a topological feature corresponds to the value of the scale parameter where it appears for the first time, and the death time of a feature corresponds to the value of the scale parameter where it dies or merges with another existing feature. Persistent homology tracks this process and records the pair of birth and death time of each topological feature in a multiset of points in $\mathbb{R}^{2}$, called persistence diagrams. One can refer to [21], [22], [37], [3] for rigorous treatment of persistent homology.
@@
Given a filtration $\{\mathcal{X}_{t}\}_{t\in\mathbb{R}}$ of a topological space $\mathcal{X}$, let, for any integer $k\geq 0$, $PH_{k}(\mathcal{X})$ denote the persistent homology of $\mathcal{X}$ that tracks the evolution of k-dimensional topological features as the scale parameter varies. Let $dgm_{k}$ denote the $k^{th}$ persistence diagram; a multiset that consists of pairs of the birth and death time of k-dimensional topological features in $PH_{k}(\mathcal{X})$. Let (b, d) correspond to the birth and death pair of a feature in $PH_{k}(\mathcal{X})$, then the persistence of the feature is defined as d - b. Features that have higher persistence are generally considered significant features, and those that correspond to lower persistence are regarded as topological noise. Another equivalent and useful representation of features is via persistence barcodes (see [12], [25]). The persistence barcode is a set of intervals where endpoints correspond to the birth and death times of a feature. A bar in the persistence barcode is the length of the interval [b, d]. Individual bars corresponding to each feature are stacked on top of each other so that the resulting structure resembles a barcode.
@@
Figure 1 depicts the persistent homology of the super-level set of a function via the persistence diagram and the persistence barcode. The leftmost figure is the graph of a non-negative function with three modes. We consider the super-level set filtration of the function to study the structures via the persistent homology. Note that at the level $t>0.20$, the super-level set of the function is empty, hence there are no non-trivial connected components or cycles to study in this case. For $t\leq 0$, the super-level set of the function equals the entire domain [-10, 10], implying that there is only one connected component in the super-level set. The interesting and non-trivial topological features occur when $t\in(0,0.20]$. Note that at $t=0.20$, the super-level set is $\{0\}$ indicating the birth of a connected component. As we decrease $t$, new elements are added to the super-level set resulting in the birth and death of topological features. The green dotted lines indicate the levels at which a new connected component is born while the black dotted lines indicate the levels at which a connected component gets merged with the existing connected components. This process of evolution of features is summarized in the persistence diagram and the persistence barcode. Let [d, b] denote an interval in the persistence barcode then b corresponds to the birth of a topological feature and d corresponds to the death of a topological feature for a super-level set filtration. For the super-level set filtration, the birth of a connected component corresponds to a local maximum, and the death of a connected component corresponds to a local minimum of the function.
@@
(Triangulable space) Given a simplicial complex $K$, the underlying space is the union of the simplices of $K$. The underlying space of $K$, when viewed as a subset of an Euclidean space, is called a polytope. We say a topological space $X$ is triangulable if there exists a finite simplicial complex with its polytope homeomorphic to $X$.

### v4

@@ Abstract
We characterize structures such as monotonicity, convexity, and modality in smooth regression curves using persistent homology. Persistent homology is a key tool in topological data analysis that detects higher-dimensional topological features such as connected components and holes (cycles or loops) in the data. In other words, persistent homology is a multiscale version of homology that characterizes sets based on the connected components and holes. We use super-level sets of functions to extract geometric features via persistent homology. In particular, we explore structures in regression curves via the persistent homology of super-level sets of a function, where the function of interest is - the first derivative of the regression function. In the course of this study, we extend an existing procedure of estimating the persistent homology for the first derivative of a regression function and establish its consistency. Moreover, as an application of the proposed methodology, we demonstrate that the persistent homology of the derivative of a function can reveal hidden structures in the function that are not visible from the persistent homology of the function itself. In particular, we characterize structures such as monotonicity, convexity, and modality, and propose a measure of statistical significance to infer these structures in practice. Finally, we conduct an empirical study to implement the proposed methodology on simulated and real data sets and compare the derived results with an existing methodology.
@@ 1 Introduction
There have been various attempts to explore the shape of the regression function from the geometric perspective, for instance; [28] proposed a rank correlation coefficient to compare the shape of regression functions. In a different line of approach, motivated by the scale-space view from computer vision literature, [6] proposed a graphical device called SiZer to explore structures, such as peaks and valleys, in regression curves. SiZer analysis has been used in comparing two or more regression curves (see, e.g., [32], [31]). [18] proposed a graphical device, based on the derivative and inverse of the derivative of the regression function, to compare the two regression functions up to shift in the horizontal and/or vertical axis. [20] compared the time-varying quantile regression curves up to horizontal shift.
@@
We analyze super-level sets of a function from a topological perspective. In particular, we take the function to be analyzed as the first derivative of the regression function and then calculate the persistent homology of its super-level sets. Since the regression function is usually not known in practice, we estimate the persistent homology using an estimate of the derivative of the regression function. In this article, we take the derivative of the Nadaraya-Watson estimator as an estimator of the derivative of the regression function. The validity of such estimators of persistent homology is due to the well-known stability theorem (see [11]). The stability theorem provides an upper bound of the bottleneck distance between true persistent homology and the estimated persistent homology in terms of the sup-norm distance between the underlying function and the corresponding estimator. Thus, if the estimator of the underlying function is uniformly convergent, then the corresponding persistent homology is consistent with respect to the bottleneck metric.
@@ 1.1 Our contribution
The persistent homology of a function does not reveal local geometric features of its own accord. It only reveals global topological features such as connected components and holes in higher dimensions. It is worth mentioning here that the use of persistent homology to study the geometric structures of 1-manifolds is mainly limited to clustering structures. The use of persistent homology to infer local geometric structures is not so straightforward, and such procedures have been less developed in the literature. In this work, we study the local geometric features such as monotonicity, convexity, and modality of smooth regression functions not by studying their persistent homology but by exploring the persistent homology of their derivative. The methodology adopted here appears analogous to the procedure proposed by [1] to estimate the persistent homology of the derivative of the regression function, though they are entirely different too. In the present work, the goal is not only to recover the persistent homology of the true function but also to infer local geometric features of the underlying function using persistent homology. In contrast, the goal in [1] is to devise a new procedure to infer the persistent homology of density and regression functions. [1] established the consistency of the proposed estimator of persistent homology for density and regression functions using their kernel estimators. In a complementary fashion, we extend those results to the first derivative of the regression function and establish the consistency of the estimator of persistent homology up to precision $5\epsilon$.
As an application of the proposed procedure, we demonstrate that it can be used to study the shape of regression functions. In particular, we characterize structures such as monotonicity, convexity, and modality in smooth regression curves and propose a significance measure of these structures observed in the data. Moreover, one can adopt the proposed procedure to compare two or more regression curves based on the structures such as monotonicity, convexity, and modality. In particular, one can perform cluster analysis for regression curves based on these structures. One of the real-life applications of our work is that, given the regression curves of population growth of different subcontinents, one can investigate the shape of the first derivative of regression functions to compare the population growth patterns over time.
@@ 1.2 Mathematical challenges
Most of the techniques in the literature first obtain an estimator of the underlying function converging uniformly to its population counterpart and then use this estimator to estimate the persistent homology. As mentioned previously, such procedures are supported by the stability results for persistent homology (see, Theorem 2.1).
@@ 1.3 Outline
The rest of the article is organized as follows. Section 2 gives the necessary topological background to elaborate the notion of persistent homology. Section 3 formulates the exact problem and the main results of the article. Section 4 presents a measure of statistical significance of observed bars. Section 5 presents characterizations of the local structures, such as monotonicity, convexity, and modality, and provides some illustrative examples to implement the proposed methodology. Section 6 contains illustrations of the proposed methodology on two real data sets. Section 7 consists of some concluding remarks. Finally, the proof of the main theoretical result is supplied in the appendix in Section 8 along with some other numerical and theoretical results.
@@ 2 Preliminaries
Persistent homology (see, e.g., [21], [25], [36]) is a prominent tool in topological data analysis (TDA) that provides a multi-scale summary of topological features present in the data. Persistent homology is a multiscale extension of homology, which is one of the refined topological invariants from algebraic topology (see, e.g., [29], [27]). Heuristically, homology characterizes sets based on the number of connected components and holes using some algebraic machinery. In what follows, we are concerned with the homology of simplicial complexes referred to as simplicial homology. Simplicial complexes are combinatorial representations of topological spaces made up of points, lines, edges, triangles, and their higher dimensional analogs. In the following subsections, we briefly describe the necessary background and definitions related to algebraic topology.
@@ 2.1 Simplicial homology
In this subsection, we first define simplicial complexes, then we describe the procedure to compute the homology of simplicial complexes.
@@
(Simplex) For any integer $k\geq 0$, a $k$-simplex is defined to be the convex hull of $k+1$ affinely independent points $\mathbb{V}=\{x_{0},x_{1},\ldots,x_{k}\}$ in $\mathbb{R}^{d},d\geq 1$. The elements of $\mathbb{V}$ are called vertices of the $k$-simplex.
In particular, points are 0-simplices, edges are 1-simplices, triangles are 2-simplices, and tetrahedra are 3-simplices.
Definition 2.2
(Faces of a simplex) Given a $k$-simplex and its vertex set $\mathbb{V}$, the faces of the $k$-simplex are defined to be the simplices spanned by the subsets of $\mathbb{V}$.
Definition 2.3
(Geometric simplicial complex) A geometric simplicial complex or a simplicial complex $\mathcal{K}$ in $\mathbb{R}^{d}$ is a collection of simplices satisfying the following conditions:
  • If $\tau$ is any face of a simplex $\sigma$ in $\mathcal{K}$, then $\tau\in\mathcal{K}$.
  • If $\tau$ and $\sigma$ are two simplices in $\mathcal{K}$, then either $\tau\cap\sigma=\emptyset$ or $\tau\cap\sigma$ is common face of both $\tau$ and $\sigma$.
Note that, when the vertex set of a geometric simplicial complex $\mathcal{K}$ is known, one can describe $\mathcal{K}$ completely and combinatorialy by the list of its simplices. This leads to the notion of an abstract simplicial complex.
Definition 2.4
(Abstract simplicial complex) Let $\mathcal{X}=\{x_{0},\ldots,x_{n}\}$ be a finite set, and $\Delta$ be a collection of non-empty subsets in $\mathcal{X}$. Then $\Delta$ is said to be an abstract simplicial complex if it satisfies the following properties:
@@
The set $\mathcal{X}$ is also called the vertex set of $\Delta$ and the elements of $\mathcal{X}$ are referred to as vertices. For an integer $k\geq 0$, any subset $\sigma\in\Delta$ with $k+1$ vertices is said to be a k-simplex or a k-dimensional simplex. Any subset $\tau$ of a k-simplex $\sigma$ with k vertices is called a face of $\sigma$. The dimension of $\Delta$ denoted as dim($\Delta$) is defined to be the maximum dimension of simplices in $\Delta$. That is, $\text{dim}(\Delta)\triangleq\text{ max }\{\text{dim}(\sigma):\sigma\in\Delta\}$.
Now, we define the following algebraic structure on simplices in $\Delta$ to define holes or cycles in different dimensions of a topological space. First, we need to assign orientations; an ordering on vertices of simplices in $\Delta$. Let $\sigma=\{x_{i_{0}},\ldots,x_{i_{k}}\}$ be a k-simplex in $\Delta$, where $i_{0},\ldots,i_{k}$ is a permutation of $0,1,\ldots,k$. We say two ordering of its vertex set is equivalent if they differ from one another by an even permutation. Thus, orderings of the vertices of $\sigma$ fall into two equivalent classes. Each of these classes is referred to as an orientation of $\sigma$. A 0-simplex has only one orientation. For any $k\geq 0$, we denote an oriented simplex as $\sigma]=[x_{i_{0}},\ldots,x_{i_{k}}$].
We denote the set of all k-simplices in $\Delta$ as $\Delta_{k}$. A k-chain on $\Delta$ is defined to be a formal linear combination of k-simplices in $\Delta_{k}$. Let $n_{k}=|\Delta_{k}|$, and $\gamma$ denotes a k-chain on $\Delta$, that is $\gamma\triangleq\sum\limits_{i=1}^{n_{k}}\alpha_{i}\sigma_{i}$, where the coefficients $\alpha_{i}$ belong to some field $\mathbb{F}$. The space of k-chains is a vector space over the field $\mathbb{F}$, and defined as
@@
Now, we need to map the simplices in $\Delta_{k}$ into a basis of $C_{k}(\mathcal{X},\mathds{F})$ such that the map preserves orientations of simplices in $\Delta_{k}$. Let $\Delta^{\pi}_{k}$ denote the set containing all the simplices in $\Delta_{k}$ in all possible orientations. Any oriented simplex $\sigma]$ in $\Delta^{\pi}_{k}$ is denoted as $x_{\pi(0)},\ldots,x_{\pi(k)}]$, where $\pi=\{\pi(0),\ldots,\pi(k)\}$ is a permutation of $0,\ldots,k$. Then, for any simplex $\sigma_{i}\in\Delta_{k}$, define the map $T_{k}:\Delta^{\pi}_{k}\xrightarrow{}C_{k}(\mathcal{X},\mathds{F})$ as follows:
@@
where $\sigma$ is a face of $\sigma_{i}$. The subspaces $B_{k}(\Delta)\triangleq\text{Im }\partial_{k+1}\text{ and }Z_{k}(\Delta)\triangleq\text{ ker }\partial_{k}$ are the set of k-boundaries and the set of k-cycles, respectively. The $k^{th}$ homology of $\Delta$ is defined to be the quotient vector space given by
@@ 2.2 Persistent homology
Let $\mathcal{X}$ be a space, then a filtration of $\mathcal{X}$ is a nested family of spaces $\{\mathcal{X}_{t}:t\in\mathbb{R}\}$ such that $\mathcal{X}=\bigcup_{t\in\mathbb{R}}\mathcal{X}_{t}$. There are two types of filtrations often used in TDA. First, filtrations built on top of a point cloud, for example, the families of Rips or $\check{C}$ech complex indexed over a set of non-negative real numbers. Second, filtrations obtained through level sets of a function $f:\mathcal{X}\xrightarrow{}\mathbb{R}$, for example, one can take $\mathcal{X}_{t}=\{x\in\mathcal{X}:f(x)\geq t\}$ to form a filtration of $\mathcal{X}$, such a filtration is referred to as a super-level set filtration of $f$. In practice, the index parameter $t$ is called the scale parameter. The homology of the spaces $\mathcal{X}_{t}$ varies with the scale parameter t. For example, as t increases, new connected components may be born, or existing connected components may disappear. Similarly, new holes or cycles are formed or filled up as t increases. Thus, as one moves through the filtration, new features are born and existing features die. The birth time of a topological feature corresponds to the value of the scale parameter where it appears for the first time, and the death time of a feature corresponds to the value of the scale parameter where it dies or merges with another existing feature. Persistent homology tracks this process and records the pair of birth and death time of each topological feature in a multiset of points in $\mathbb{R}^{2}$, called persistence diagrams. One can refer to [21], [22], [36], [3] for rigorous treatment of persistent homology.
@@
Given a filtration $\{\mathcal{X}_{t}\}_{t\in\mathbb{R}}$ of a topological space $\mathcal{X}$, let, for any integer $k\geq 0$, $PH_{k}(\mathcal{X})$ denote the persistent homology of $\mathcal{X}$ that tracks the evolution of k-dimensional topological features as the scale parameter varies. Let $dgm_{k}$ denote the $k^{th}$ persistence diagram; a multiset that consists of pairs of the birth and death time of k-dimensional topological features in $PH_{k}(\mathcal{X})$. Let (b, d) correspond to the birth and death pair of a feature in $PH_{k}(\mathcal{X})$, then the persistence of the feature is defined as d - b. Features that have higher persistence are generally considered significant features, and those that correspond to lower persistence are regarded as topological noise. Another equivalent and useful representation of features is via persistence barcodes (see [12], [25]). The persistence barcode is a set of intervals where endpoints correspond to the birth and death time of a feature. A bar in the persistence barcode is the length of the interval [b, d]. Individual bars corresponding to each feature are stacked on top of each other so that the resulting structure resembles a barcode.
@@
Figure 1 depicts the persistent homology of the super-level set of a function via the persistence diagram and the persistence barcode. The leftmost figure is the graph of a non-negative function with three modes. We consider the super-level set filtration of the function to study the structures via the persistent homology. Note that at the level $t>0.20$, the super-level set of the function is empty, hence there are no non-trivial connected components or cycles to study in this case. For $t\leq 0$, the super-level set of the function equals the entire domain [-10, 10], implying that there is only one connected component in the super-level set. The interesting and non-trivial topological features occur when $t\in(0,0.20]$. Note that at $t=0.20$, the super-level set is $\{0\}$ indicating the birth of a connected component. As we decrease $t$, new elements are added to the super-level set, resulting in the birth and death of topological features. The green dotted lines indicate the levels at which a new connected component is born, while the black dotted lines indicate the levels at which a connected component gets merged with the existing connected components. This process of evolution of features is summarized in the persistence diagram and the persistence barcode. Let [d, b] denote an interval in the persistence barcode; then b corresponds to the birth of a topological feature and d corresponds to the death of a topological feature for a super-level set filtration. For the super-level set filtration, the birth of a connected component corresponds to a local maximum, and the death of a connected component corresponds to a local minimum of the function.
@@
(Triangulable space) Given a geometric simplicial complex $\mathcal{K}$, the underlying space of $\mathcal{K}$ is the union of the simplices of $\mathcal{K}$. The underlying space of $\mathcal{K}$, when viewed as a subset of an Euclidean space, is called a polytope. We say a topological space $X$ is triangulable if there exists a finite simplicial complex with its polytope homeomorphic to $X$.